What is the molarity of NO?3 in each solution?

0.150 M KNO3.

0.220 M Ca(NO3)2

0.370 M Al(NO3)3.
Answer #1

KNO3................> K+ + NO3-

KNO3 produce one NO3- ion.

so molarity of NO?3 in this solution = 0.150 M

0.220 M Ca(NO3)2:

Ca(NO3)2......................> Ca2+ + 2 NO3-

Ca(NO3)2 produce two NO3- ion.

so molarity of NO?3 = 2 * 0.220 = 0.440 M

0.370 M Al(NO3)3.:

Al(NO3)3 .............> Al3+ + 3 NO3-

Al(NO3)3 produce three NO3- ions.

molarity of NO?3 = 3 * 0.370 = 1.11 M